OpenStudy (anonymous):

how many real roots does the equation x^7 + x have

OpenStudy (allank):

Hmm...that is an expression. You need to add "=sumthing" to make it an equation. Maybe it's x^7+x=0 ?

OpenStudy (anonymous):

oh sorry i forgot to add that part

OpenStudy (allank):

No prob. Let's work it together. First we factor out x: x(x^6+1)=0 okay?

OpenStudy (anonymous):

okay

OpenStudy (allank):

So that means that either x = 0 (first real root) or x^6+1=0 ok?

OpenStudy (anonymous):

i see, so we just solve it by factoring? no need to use the rule of signs or anything fancy?

OpenStudy (allank):

We will do something a bit fancy with x^6+1=0 So that means that x^6=-1 Now, there is no real value for x that can be raised to power 6, to get a value of -1. That means that there is only one real root for the original equation: x=0

OpenStudy (anonymous):

okay, thanks for the help

OpenStudy (allank):

You're welcome :)